# Trivialization of the gravitational Green-Schwarz mechanism in the non-relativistic limit of string theory

### v1

@@ Abstract
We show that the gravitational Green–Schwarz anomaly cancellation mechanism becomes trivial in the non-relativistic (NR) limit of ten-dimensional heterotic supergravity with four-derivative corrections. We perform a NR expansion of the Kalb–Ramond field and identify the finite Green–Schwarz transformation in this limit, which can be interpreted as a non-covariant $SO(8)$ transformation. We then construct an explicit field redefinition such that the redefined two-form is invariant under this symmetry. This result is compared with the previously reported trivialization of the gauge Green–Schwarz mechanism under the same limit. Both field redefinitions can be implemented simultaneously, and the associated Chern–Simons terms are exact, arising directly from the redefinition structure, and leading to a trivial Bianchi identity. These results imply that anomaly cancellation becomes automatic in the NR regime, extending the consistent formulations of NR heterotic string theories and giving rise to a richer class of topological heterotic backgrounds with simpler thermodynamical properties such as the Wald entropy and the first law.
@@ I Introduction
In this work, we demonstrate that the Green-Schwarz mechanism becomes trivial in the NR limit of heterotic supergravity. More precisely, we show that up to four-derivative order (i.e., to first order in $\alpha^{\prime}$), the nontrivial gravitational Green-Schwarz transformation of the $B$-field can be entirely removed by field redefinitions. As a consequence, the redefined two-form $\tilde{b}_{\mu\nu}$ becomes manifestly invariant under $SO(8)$ transformations. Moreover, the modified three-form field strength $\bar{h}_{\mu\nu\rho}$ becomes exact, and its Bianchi identity is trivially satisfied. These results, together with the previously reported gauge Green-Schwarz trivialization Lescano2025 , have profound implications: the anomaly cancellation in heterotic string theory no longer imposes constraints on the gauge group or background topology. The need for anomaly inflow is eliminated, and the associated cohomological conditions are no longer required. This leads to a dramatic relaxation of consistency constraints and allows for compactifications and gauge sectors that would be anomalous in the relativistic theory. Similarly, the Wald entropy and the first law of thermodynamics are simplified under this limit, thanks to the simplification of the Bianchi identity for the three-form curvature. These findings open the door to a broader landscape of consistent NR heterotic theories, including those with unconventional gauge groups, torsional geometries, or non-Kähler compactifications upon considering the NR limit. Furthermore, it provides a new perspective on the role of anomalies in string theory, suggesting that non-Lorentzian limits can fundamentally alter their geometric and topological origin. The four-derivative corrections to the universal action principle of the different formulations of heterotic string theory were historically computed considering three- and four-point scattering amplitudes for the massless states GrossSloan -CN . This method is based on the study of the different types of string interactions through the S-matrix to construct an effective Lagrangian, originally computed by Metsaev and TseytlinMetsaevTseytlin ,
@@ II.2 Field redefinitions and Green-Schwarz trivialization
$\delta_{\lambda}b_{\mu\nu}=-\frac{\alpha^{\prime}_{NR}}{4}\partial_{\mu}\lambda^{a^{\prime}b^{\prime}}\tau_{\nu]}{}^{d}\eta_{de}\tau_{\rho\sigma}{}^{e}e^{\rho}{}_{a^{\prime}}e^{\sigma}{}_{b^{\prime}}\,,$ (14)
@@
$C_{\mu\nu\rho}=-\frac{\alpha^{\prime}_{NR}}{4}\partial_{[\mu}\Big{(}w_{\nu}{}^{a^{\prime}b^{\prime}}\tau_{\rho]}{}^{d}\eta_{de}e^{\epsilon}{}_{a^{\prime}}e^{\sigma}{}_{b^{\prime}}\tau_{\epsilon\sigma}{}^{e}\Big{)}\,.$ (20)
@@ III Implications
The trivialization of the gravitational and gauge Green-Schwarz mechanism in the NR limit of heterotic supergravity leads to a profound restructuring of the constraints of heterotic string theory. We now outline the main consequences.
@@ III.1 Anomaly cancellation
In relativistic heterotic string theory, one-loop gauge and gravitational anomalies are canceled by a nontrivial transformation of the Kalb-Ramond $B$-field, via its coupling to Chern-Simons terms GS . In the NR limit, the Green-Schwarz transformations are absorbed through field redefinitions as showed in the previous section, and therefore the redefined $\tilde{b}_{\mu\nu}$ field is invariant under both local Lorentz and gauge transformations. Consequently, the need for anomaly inflow is eliminated: anomalies no longer appear in the effective action, up to covariant field redefinitions. This trivialization suggests that anomaly cancellation is no longer an external consistency condition, in the NR limit of heterotic string theory. In particular, it implies that the anomaly polynomial becomes cohomologically trivial at this order.
@@ III.3 Modified topological constraints in compactifications
identically, without requiring a nontrivial cancellation between $\mathrm{tr}(F\wedge F)$ and $\mathrm{tr}(R\wedge R)$. As a result, the anomaly inflow condition becomes structurally trivial, and no cohomological constraint on gauge or tangent bundles remains. This opens up the possibility of previously forbidden compactifications including those with non-matching Chern classes, non-Kähler internal geometries, or background torsion in the NR limit.
@@ III.4 Wald entropy and the first law in non-relativistic heterotic backgrounds
The analysis of black hole thermodynamics in the presence of higher-derivative corrections has traditionally relied on Wald’s formalism Wald . For heterotic supergravity at order $\alpha^{\prime}$, the relevant Noether charges require careful treatment due to the Green-Schwarz couplings between the Kalb-Ramond field and the gauge and gravitational Chern-Simons termsOrtin . In the NR limit, the results of this letter predict a dramatic simplification of this structure. Since both the gauge and Lorentz Green-Schwarz mechanisms are trivialized, the redefined Kalb-Ramond field transforms covariantly, and no Chern-Simons modifications are induced. As a result, the Bianchi identity for the corrected three-form curvature reduces to a standard closure condition in the NR limit, as in Eq. (29). This implies that the horizontal Kalb-Ramond momentum map $P_{k}$ becomes simpler, relaxing its role in the entropy computation (with a similar argument we expect a relaxation in the vertical Yang-Mills momentum map). The trivialization of the Green-Schwarz mechanism in the NR limit of heterotic string theory thus enables a simpler derivation of the first law of black hole mechanics, in which all thermodynamic potentials and conserved charges are computed from manifestly gauge- and Lorentz-invariant expressions, up to four-derivative contributions.
@@ III.5 Connection to duality-invariant frameworks
The NR limit of heterotic string theory aligns naturally with duality-invariant formulations such as Double Field Theory (DFT), where generalized geometry encodes both geometric and gauge degrees of freedom EandD . The four-derivative extension of DFT Tduality Hohm involves a generalized vielbein formalism DFTframe . It is therefore not clear whether the NR limit produces a finite four-derivative Lagrangian considering the flux formulation of DFT, compatible with a generalization of the Green-Schwarz mechanism Odd . This motivates a complete NR extension of heterotic DFT, incorporating supersymmetry and $\alpha^{\prime}$ corrections beyond leading order, extending the results of LNR for the relativistic case.
@@ IV Conclusion
We have proved that the Green-Schwarz mechanism of heterotic string theory can be trivialized in the NR limit up to four-derivative terms. We constructed the field redefinition of the b-field that makes this possible and, by construction, the Bianchi identity for the 3-form curvature does not include $\alpha^{\prime}$-corrections. Our results establish a new paradigm for anomaly cancellation in the NR limit of heterotic string theory with several implications and open directions for future research in the long term. The immediate continuation of this work includes:
  • Investigating the role of supersymmetry when higher-derivative terms are considered, extending the seminal work in Bsusy and exploring further trivializations of the supersymmetric transformations at the NR supergravity level Bdroo .
  • Understanding the embedding of the duality invariant construction of LNR for the higher-derivative structure of heterotic DFT in the NR limit, including the construction of the full four-derivative heterotic finite action (including fermions) in the NR limit.

### v2

# Trivialization of the gravitational Green-Schwarz transformation in the non-relativistic limit of string theory

@@ Abstract
We show that the gravitational Green–Schwarz (GS) transformation becomes trivial in the non-relativistic (NR) limit of ten-dimensional heterotic supergravity with four-derivative corrections. This constitutes an important step towards establishing the trivialization of the GS mechanism in this limit. In this work, we perform a NR expansion of the Kalb–Ramond field and identify the finite Green–Schwarz transformation in this limit, which can be interpreted as a non-covariant $SO(8)$ transformation. We then construct an explicit field redefinition such that the redefined two-form is invariant under this symmetry. This result is compared with the previously reported trivialization of the gauge GS mechanism under the same limit. Both field redefinitions can be implemented simultaneously, and the associated Chern–Simons terms are exact, arising directly from the redefinition structure, and leading to a trivial Bianchi identity. These results support the expectation that anomaly cancellation becomes automatic in the NR regime, and therefore we discuss their potential implications.
@@ I Introduction
While a previous study Lescano2025 analyzed the gauge Green–Schwarz mechanism in the NR limit—focusing on the behavior of the gauge sector and its associated two-derivative action, the present Letter addresses a different and independent problem: the gravitational Green–Schwarz mechanism and its four-derivative $\alpha^{\prime}$-corrections. The two mechanisms involve distinct symmetries and Chern–Simons structures: the gauge mechanism concerns the non-Abelian field strength and its coupling to the $B$-field, while the gravitational mechanism involves the Lorentz connection and curvature.
In this work, we demonstrate that the gravitational Green-Schwarz transformation of the $b$-field can be entirely removed by field redefinitions. As a consequence, the redefined two-form $\tilde{b}_{\mu\nu}$ becomes manifestly invariant under $SO(8)$ transformations. Moreover, the modified three-form field strength $\bar{h}_{\mu\nu\rho}$ becomes exact, and its Bianchi identity is trivially satisfied. These results, together with the previously reported gauge Green-Schwarz trivialization Lescano2025 , have profound implications: the anomaly cancellation in heterotic string theory may no longer impose constraints on the gauge group or background topology. The need for anomaly inflow could be eliminated, and the associated cohomological conditions would no longer be required. While in this work we pave the way towards the full trivialization of the Green-Schwarz mechanism, a complete cancellation would lead to a dramatic relaxation of consistency constraints and allows for compactifications and gauge sectors that would be anomalous in the relativistic theory. Similarly, the Wald entropy and the first law of thermodynamics are simplified under this limit, thanks to the simplification of the Bianchi identity for the three-form curvature. These findings open the door to a broader landscape of consistent NR heterotic theories in the long term, including those with unconventional gauge groups, torsional geometries, or non-Kähler compactifications upon considering the NR limit. Furthermore, if the full Green-Schwarz mechanism would be trivial under this limit, this could provide a new perspective on the role of anomalies in string theory, suggesting that non-Lorentzian limits can fundamentally alter their geometric and topological origin. The four-derivative corrections to the universal action principle of the different formulations of heterotic string theory were historically computed considering three- and four-point scattering amplitudes for the massless states GrossSloan -CN . This method is based on the study of the different types of string interactions through the S-matrix to construct an effective Lagrangian, originally computed by Metsaev and TseytlinMetsaevTseytlin ,
@@ II.2 Field redefinitions and Green-Schwarz trivialization
$\delta_{\lambda}b_{\mu\nu}=-\frac{\alpha^{\prime}_{NR}}{4}\partial_{[\mu}\lambda^{a^{\prime}b^{\prime}}\tau_{\nu]}{}^{d}\eta_{de}\tau_{\rho\sigma}{}^{e}e^{\rho}{}_{a^{\prime}}e^{\sigma}{}_{b^{\prime}}\,$, (14)
@@
[…]
@@ III Implications
A full trivialization of the gravitational and gauge Green-Schwarz mechanism in the NR limit of heterotic supergravity would lead to a profound restructuring of the constraints of heterotic string theory. Motivated by the initial results in this paper, where we trivialize the gravitational Green-Schwarz transformation, we now outline the main consequences of the full trivialization of the Green-Schwarz mechanism.
@@ III.1 Anomaly cancellation
In relativistic heterotic string theory, one-loop gauge and gravitational anomalies are canceled by a nontrivial transformation of the Kalb-Ramond $B$-field, via its coupling to Chern-Simons terms GS . In the NR limit, the Green-Schwarz transformation of the b-field is absorbed through a field redefinition, as showed in the previous section, and therefore the redefined $\tilde{b}_{\mu\nu}$ field is invariant under both local Lorentz and gauge transformations. Consequently, it is expected that all the heterotic anomalies could be eliminated through field redefinitions and no longer appear in the effective action  [^2] The analysis of the anomalies still requires the implementation of covariant field redefinitions, which are not straightforward to implement. The best strategy is to study the action at higher-derivative orders, and guide the cancellation with the field redefinitions used to make the action finite. The author thanks his anonymous referee for this observation.. The trivialization of this work therefore suggests that anomaly cancellation may no longer be an external consistency condition, in the NR limit of heterotic string theory. In particular, it could potentially imply that the anomaly polynomial becomes cohomologically trivial at this order.
@@ III.3 Modified topological constraints in compactifications
identically, without requiring a nontrivial cancellation between $\mathrm{tr}(F\wedge F)$ and $\mathrm{tr}(R\wedge R)$. As a result, the anomaly inflow condition may become structurally trivial, and no cohomological constraint on gauge or tangent bundles may remain. The results of this paper, together with a complete prove of the Green-Schwarz mechanism, open up the possibility of previously forbidden compactifications including those with non-matching Chern classes, non-Kähler internal geometries, or background torsion in the NR limit.
@@ III.4 Wald entropy and the first law in non-relativistic heterotic backgrounds
The analysis of black hole thermodynamics in the presence of higher-derivative corrections has traditionally relied on Wald’s formalism Wald . For heterotic supergravity at order $\alpha^{\prime}$, the relevant Noether charges require careful treatment due to the Green-Schwarz couplings between the Kalb-Ramond field and the gauge and gravitational Chern-Simons termsOrtin . In the NR limit, the results of this letter predict a dramatic simplification of this structure. Since both the gauge and Lorentz Green-Schwarz mechanisms are trivialized, the redefined Kalb-Ramond field transforms covariantly, and no Chern-Simons modifications are induced. As a result, the Bianchi identity for the corrected three-form curvature reduces to a standard closure condition in the NR limit, as in Eq. (29). This implies that the horizontal Kalb-Ramond momentum map $P_{k}$ becomes simpler, relaxing its role in the entropy computation (with a similar argument we expect a relaxation in the vertical Yang-Mills momentum map). The trivialization of the Green-Schwarz transformation in the NR limit of heterotic string theory thus enables a simpler derivation of the first law of thermodynamics, in which all thermodynamic potentials and conserved charges are computed from manifestly gauge- and Lorentz-invariant expressions, up to four-derivative contributions.
@@ III.5 Connection to duality-invariant frameworks
The NR limit of heterotic string theory aligns naturally with duality-invariant formulations such as Double Field Theory (DFT) DFT1 -DFT2 , where generalized geometry encodes both geometric and gauge degrees of freedom EandD . The four-derivative extension of DFT Tduality Hohm involves a generalized vielbein formalism DFTframe . It is therefore not clear whether the NR limit produces a finite four-derivative Lagrangian considering the flux formulation of DFT, compatible with a generalization of the Green-Schwarz mechanism Odd . This motivates a complete NR extension of heterotic DFT, incorporating supersymmetry and $\alpha^{\prime}$ corrections beyond leading order, extending the results of LNR in the NR limit.
@@ IV Conclusion
We have proved that the gravitational Green-Schwarz transformation of heterotic supergravity can be trivialized in the NR limit up to four-derivative terms. We constructed the field redefinition of the b-field that makes this possible and, by construction, the Bianchi identity for the 3-form curvature does not include $\alpha^{\prime}$-corrections. Our results open the door to a new paradigm for anomaly cancellation in the NR limit of heterotic string theory with several implications if the full Green-Schwarz mechanism can be proved trivial. The immediate continuation of this work includes:
  • Studying the role of supersymmetry when higher-derivative terms are considered, extending the seminal work in Bsusy and exploring further trivializations of the supersymmetric transformations at the NR supergravity level Bdroo .
  • Understanding the embedding of the duality invariant construction of LNR for the higher-derivative structure of heterotic DFT in the NR limit, including the construction of the full four-derivative heterotic finite action (including fermions) in the NR limit. Initial progress was already done in LescanoExtra using the Bergshoeff-de Roo identification.
@@
  • Extending the analysis to six- and eight-derivatives order, towards a complete proof of the trivialization of the Green-Schwarz mechanism under the NR limit.